# Learning with less: label-efficient land cover classification at very high spatial resolution using self-supervised deep learning

### v1

@@ 5 Discussion
Core to the success of our approach is the use of self-supervised learning to pre-train an image encoder using large amounts of unannotated NAIP imagery prior to fine-tuning on the target land cover classification task. We achieved high land cover classification accuracy with only 1,000 labeled training samples, which is a small fraction of the amount of labeled data typically used to train deep learning models for similar tasks. For example, the ISPRS Potsdam and Vaihingen datasets can be used to create 23,750 and 8,873 256256 pixel training patches, respectively [40, 41]. Additionally, these datasets only cover small urban areas, while our work concerns a substantially larger extent with multiple diverse ecoregions. The results of our experiment provide strong evidence to support this claim, and we applied this methodology to produce a resolution land cover map for the entire state of Mississippi. We note that for linear probing, the embeddings produced by the encoder pre-trained using BYOL alone performed the best, slightly outperforming the BYOL model initialized with ImageNet weights and substantially outperforming the ImageNet pre-trained model. One observation from the quantitative assessment of the model is the discrepancy between the performance of the pre-trained models when used as fixed feature extractors for linear probing compared to when they are fine-tuned as part of an end-to-end trainable semantic segmentation architecture. When only working with fixed embeddings, the BYOL model with random initialization outperformed the BYOL model initialized with ImageNet weights, which in turn outperformed the ImageNet pre-trained model. This suggests that the embeddings produced by the BYOL model with random initialization are somewhat linearly separable with respect to the land cover classes of interest, even in the absence of any supervised fine-tuning. This is a promising result, as it contributes to the growing body of evidence that self-supervised learning methods can be just as effective (if not more so) than traditional supervised pre-training methods for remote sensing applications [106, 107, 108]. These results also provided a quantitative demonstration of the domain gap between the natural, general-purpose images found in ImageNet and remote sensing imagery, and the benefits of pre-training on in-domain data, as echoed by findings in other recent works [109]. However, when fine-tuning the entire model as part of a semantic segmentation architecture, the BYOL model initialized with ImageNet weights consistently outperformed both the BYOL model with random initialization and the ImageNet pre-trained model across all architectures and training dataset sizes. This is a critical observation, as it indicates that the broader, more general image features learned from ImageNet are potentially relevant to the land cover classification task when a more complex decoder network is used to learn task-specific features during fine-tuning, suggesting that in spite of the apparent domain gap, ImageNet pre-training is still worthwhile for remote sensing applications.
@@
We note that while our results demonstrate the effectiveness of BYOL for pre-training an image encoder for land cover classification at VHSR, other self-supervised learning methods may also be effective for this task. One notable restriction of BYOL and similar transformation-based self-supervised learning methods is that they tend to rely on image augmentation routines that are specifically designed for three-channel RGB imagery, which can limit their applicability to multispectral or hyperspectral remote sensing imagery that contains more than three bands. For example, a central component of BYOL’s augmentation pipeline involves adjusting the brightness, contrast, saturation, and hue of an input image. The saturation and hue adjustment transformations are explicitly defined for three-channel imagery as a mapping from the RGB color space to the HSV color space, and thus cannot be directly applied to imagery with more than three bands. For remote sensing applications, this is a significant limitation of transformation-based self-supervised learning methods that rely on image augmentations to learn transformation-invariant image representations, restricting their applicability to only 3-band imagery. While we achieved strong results using BYOL on 3-band color-infrared NAIP imagery, 4 or more bands may be necessary to effectively utilize spectral information for other remote sensing applications, such as vegetation mapping, crop classification, or soil property estimation. Future work should continue to explore the implications of self-supervised learning for remote sensing applications under real-world constraints beyond academic benchmarks and compute-intensive techniques. Further, for label-scarce scenarios, little information on optimal sampling strategies for patch-based dense classification tasks is currently available in the literature, as opposed to more well-studied pixel-based sampling strategies for traditional machine learning classifiers [117, 75, 6].

### v2

@@ 5 Discussion
Core to the success of our approach is the use of self-supervised learning to pre-train an image encoder using large amounts of unannotated NAIP imagery prior to fine-tuning on the target land cover classification task. We achieved high land cover classification accuracy with only 1,000 labeled training samples, which is a small fraction of the amount of labeled data typically used to train deep learning models for similar tasks. For example, the ISPRS Potsdam and Vaihingen datasets can be used to create 23,750 and 8,873 256256 pixel training patches, respectively [53, 54]. This corresponds to an approximately 96% and 89% reduction in the quantity of labeled data required compared to these benchmarks, representing a substantial decrease in the manual resources necessary for annotation. Additionally, these datasets only cover small urban areas, while our work concerns a substantially larger extent with multiple diverse ecoregions. The results of our experiment provide strong evidence to support this claim, and we applied this methodology to produce a resolution land cover map for the entire state of Mississippi. We note that for linear probing, the embeddings produced by the encoder pre-trained using BYOL alone performed the best, slightly outperforming the BYOL model initialized with ImageNet weights and substantially outperforming the ImageNet pre-trained model. One observation from the quantitative assessment of the model is the discrepancy between the performance of the pre-trained models when used as fixed feature extractors for linear probing compared to when they are fine-tuned as part of an end-to-end trainable semantic segmentation architecture. When only working with fixed embeddings, the BYOL model with random initialization outperformed the BYOL model initialized with ImageNet weights, which in turn outperformed the ImageNet pre-trained model. This suggests that the embeddings produced by the BYOL model with random initialization are somewhat linearly separable with respect to the land cover classes of interest, even in the absence of any supervised fine-tuning. This is a promising result, as it contributes to the growing body of evidence that self-supervised learning methods can be just as effective (if not more so) than traditional supervised pre-training methods for remote sensing applications [63, 5, 116]. These results also provided a quantitative demonstration of the domain gap between the natural, general-purpose images found in ImageNet and remote sensing imagery, and the benefits of pre-training on in-domain data, as echoed by findings in other recent works [93]. However, when fine-tuning the entire model as part of a semantic segmentation architecture, the BYOL model initialized with ImageNet weights consistently outperformed both the BYOL model with random initialization and the ImageNet pre-trained model across all architectures and training dataset sizes. This is a critical observation, as it indicates that the broader, more general image features learned from ImageNet are potentially relevant to the land cover classification task when a more complex decoder network is used to learn task-specific features during fine-tuning, suggesting that in spite of the apparent domain gap, ImageNet pre-training is still relevant for remote sensing applications.
We found that across all cases, the BYOL consistently outperformed MoCoV2 in both linear probing and fine-tuning performance. This is notable as BYOL is ultimately a simpler approach to pre-training compared to MoCoV2, foregoing requirements for large numbers of negative samples required by the contrastive loss function and requiring fewer hyper-parameters to tune. For example, beyond the standard set of hyper-parameters for model training, MoCoV2 performance tends to vary under different values for momentum (parameter in Equation 2), temperature (parameter in Equation 3), queue size, and projection head embedding dimension, while BYOL only requires tuning of the momentum parameter and projection head embedding dimension. For all experiments, we used parameters discovered by the original authors of the respective works determined to yield good results; however, given the domain shift between general purpose images and aerial imagery, it is possible that tuning these hyperparameters may yield additional performance. However, tuning these is computationally expensive due to the large amount of training data required and long training schedules. Additionally, the contrastive loss may be ill-posed for this task due to the homogeneity of the pre-training dataset. For example, Figure S2 shows that in the pre-training data, many separate samples share very similar characteristics due to the fact that a large portion of Mississippi’s landscape is dominated by forested land cover or cultivated crops. When using a contrastive loss, the model may be penalized for outputting similar embeddings for negative pairs even when those negative pairs share similar semantic content (i.e., “false negatives”). For example, even though two tiles consist completely of forested area, they are still treated as negative pairs for the purpose of optimization. Increasing the temperature parameter to smooth the optimization landscape or reducing the queue size may help alleviate these issues, but this is a well-recognized fundamental flaw in applying contrastive learning to homogeneous datasets [52, 19, 149]. BYOL’s self-distillation framework avoids this scenario entirely by incorporating asymmetry in the pre-training in order to give the online encoder a robust training signal without the need for negative samples.
@@
In terms of architectures, we found that the U-Net models tended to outperform all other architectures tested, despite being among the simpler frameworks for semantic segmentation. One potential reason for this is the use of a symmetric decoder framework with skip connections instead of a shallow decoder network (as used in DeepLabV3+ and FCN). This can also be observed based on the performance of the Attention U-Net and UPerNet architectures, which follow similar symmetric encoder-decoder architectures. However, these models integrate complex sub-networks in an attempt to improve performance, such as attention gates and pyramid pooling. While these sub-networks may facilitate better multi-scale feature extraction or spatial feature localization when sufficient data is available for training, under limited data scenarios they appear to hinder the performance of the full networks [10]. On the other hand, the U-Net decoder consists entirely of 33 convolutions, batch normalization, and bilinear upsampling layers: this has the effect of making it straightforward to fit with few training samples, but reduces the effective receptive field of the network, making incorporating long-range context into classifications difficult. Section S2 in the supplementary materials includes a short discussion on how this impacts the classification of large buildings, leading to confusion between the impervious surfaces and impervious structures classes.
@@
We note that while our results demonstrate the effectiveness of BYOL for pre-training an image encoder for land cover classification at VHSR, other self-supervised learning methods may also be effective for this task. One notable restriction of BYOL, MoCoV2, and similar transformation-based self-supervised learning methods is that they tend to rely on image augmentation routines that are specifically designed for three-channel RGB imagery, which can limit their applicability to multispectral or hyperspectral remote sensing imagery that contains more than three bands. For example, a central component of BYOL’s augmentation pipeline involves adjusting the brightness, contrast, saturation, and hue of an input image. The saturation and hue adjustment transformations are explicitly defined for three-channel imagery as a mapping from the RGB color space to the HSV color space, and thus cannot be directly applied to imagery with more than three bands. For remote sensing applications, this is a significant limitation of transformation-based self-supervised learning methods that rely on image augmentations to learn transformation-invariant image representations, restricting their applicability to only 3-band imagery. While we achieved strong results using BYOL on 3-band CIR NAIP imagery, 4 or more bands may be necessary to effectively utilize spectral information for other remote sensing applications, such as vegetation mapping, crop classification, or soil property estimation. In the case of land cover classification, we found that CIR composites are sufficient for robust land cover classification tasks, especially given that spectral reflectance in the blue band is unreliable due to atmospheric scattering [104, 58, 124], and historically has provided little utility for other land cover/land use classification tasks [2, 118]. Still, our model struggles to differentiate between impervious surfaces and barren land (see Table 3 and sample (b) in Figure S1 in the supplementary materials), and given the utility of the blue band for soil mapping [73, 105], the inclusion of this band could increase the spectral separability of the two classes. Future work should continue to explore the implications of self-supervised learning for multispectral remote sensing applications under real-world constraints beyond academic benchmarks and compute-intensive techniques.
Further, for label-scarce scenarios, little information on optimal sampling strategies for patch-based dense classification tasks is currently available in the literature, as opposed to more well-studied pixel-based sampling strategies for traditional machine learning classifiers [13, 155, 112]. Our approach relied on using existing coarse-resolution land cover data to select a representative set of training samples, but this approach may introduce bias into the training dataset given that the NLCD data is derived from another classification algorithm, not a human annotator, and the resolution is too coarse to capture fine details in land cover, such as small roads or barren clearings (see Figure 11). Subsequently, the feature vectors derived from this data and used for stratification may not be representative of high-frequency spatial land cover patterns, introducing bias into the dataset. Examining the ratio between number of predicted samples and number of actual samples for each class in Table 3 gives an insight into specific classes that our model is biased towards or against. For example, in the 2023 assessment, the model predicted 1,886 pixels as cultivated crops, while the actual number of pixels corresponding cultivated crops is only 1,301. On the other hand, the model predicted only 1,142 pixels as belonging to the barren land class, while 1,644 pixels actually belonged to the barren land class. Relying on such a data product may bias the model towards classes that are well-represented or well-classified according to the coarse resolution multitemporal datasets used to produce the land cover dataset, and against classes that are too small to be apparent in the coarse data (e.g., small barren clearings) or rely on phenological characteristics (e.g., crop fields that have been harvested prior to NAIP imagery acquisition). Given the absence of literature on best practices for designing stratification protocols for sampling training data for patch-based classifiers, further analysis is needed to fully understand the implications of such approaches.
@@ Acknowledgments
Supplementary material
S1 Model performance breakdown
| Model | Pre-train Method | Overall Acc. (%) | User’s Acc. (%) | Producer’s Acc. (%) | F1 Score (%) | IoU (%) | Cohen’s | |
| Linear Probe | ImageNet Only | 250 | ||||||
| 500 | ||||||||
| 750 | ||||||||
| MoCoV2 Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + MoCoV2 | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| BYOL Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + BYOL | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| FCN | ImageNet Only | 250 | ||||||
| 500 | ||||||||
| 750 | ||||||||
| MoCoV2 Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + MoCoV2 | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| BYOL Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + BYOL | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| PAN | ImageNet Only | 250 | ||||||
| 500 | ||||||||
| 750 | ||||||||
| MoCoV2 Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + MoCoV2 | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| BYOL Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + BYOL | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| DeepLabV3+ | ImageNet Only | 250 | ||||||
| 500 | ||||||||
| 750 | ||||||||
| MoCoV2 Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + MoCoV2 | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| BYOL Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + BYOL | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| UPerNet | ImageNet Only | 250 | ||||||
| 500 | ||||||||
| 750 | ||||||||
| MoCoV2 Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + MoCoV2 | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| BYOL Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + BYOL | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| Attention U-Net | ImageNet Only | 250 | ||||||
| 500 | ||||||||
| 750 | ||||||||
| MoCoV2 Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + MoCoV2 | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| BYOL Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + BYOL | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| U-Net | ImageNet Only | 250 | ||||||
| 500 | ||||||||
| 750 | ||||||||
| MoCoV2 Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + MoCoV2 | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| BYOL Only | 250 | |||||||
| 500 | ||||||||
| 750 | ||||||||
| ImageNet + BYOL | 250 | |||||||
| 500 | ||||||||
| 750 |
S2 Assessment of model misclassifications
Examining the spatial distribution of errors in Figure S1 reveals that errors are most heavily concentrated in the Mississippi Delta region, where agricultural activities are the predominant land use and a large amount of area is dedicated to cultivated crops. Table 4 shows that our model has a high producer’s accuracy for cultivated crops (95.54%), but poor user’s accuracy (65.91%), indicating that while the model can accurately classify actual crop fields, it has a propensity to misclassify herbaceous land cover types as cultivated crops. An example of such a misclassification is shown in samples (d) and (e) in Figure S1, where portions of herbaceous fields are misclassified as cropland. The class definitions in Table 1 describes how both strong near-infrared reflectance (for crops near peak growth) and the presence of row patterns are potential indicators of cropland. However, for crops where the spacing between rows is not visible in the imagery, or in cases where the imagery is acquired when the crop is not near peak growth, classification can be more complex. To account for this, the model overclassified cultivated crop regions, leading to highly accurate classifications of regions that are actually cropland, but a tendency to commit errors of commission when classifying other vegetated land cover classes (e.g., herbaceous).
Unlike cultivated crops, the barren land class is prone to under-classification, with a moderately strong user’s accuracy of 82.66% but a weak producer’s accuracy of 57.42%. Barren land regions are typically characterized by a lack of any texture and low near-infrared reflectance, indicating a total lack of vegetation. However, for regions with mixed herbaceous/barren land cover types, the distinction between the two can be difficult to delineate precisely. Sample (d) in Figure S1 shows an example of this: while the herbaceous regions show some texture, it is difficult to visually distinguish where the herbaceous land cover ends and where the barren land begins. Further, impervious surfaces are very similar to barren land cover types in terms of both spatial and spectral features. Sample (b) in Figure S1 demonstrates both errors of commission and omission with regards to the misclassification of barren land: the middle of the sample shows a small stream with barren banks and deposits, but narrow portions are misclassified as impervious surfaces, while the impervious road in the upper part of the image is partially misclassified as barren land. In the absence of additional spectral information or ancillary road data, separating the two classes based on the color-infrared imagery alone is difficult.
From a numerical standpoint, classification of impervious structures (i.e., buildings) is moderately high with similar user’s and producer’s accuracy (76.19% and 75.47%, respectively). Most misclassification of buildings appeared to arise from lack of nearby spatial features that aid in distinguishing impervious surfaces from impervious structures, such as shadows (indicating elevation) or hard edges/corners. Sample (c) in Figure S1 demonstrates the effect this has on classifications in dense urban landscapes, where large buildings have a tendency to be misclassified as impervious surfaces, while smaller buildings were classified more accurately. This could be a limitation of the use of CNNs, which have limited receptive fields and struggle to incorporate long-range contextual features into classifications [80, 74].
Systematic errors arising from data quality issues are also present in the final product. One such issue that is prevalent in rivers/waterways is the presence of shipping barges and other watercraft that cause misclassifications in open water areas. Such watercraft are mostly present in the Mississippi River due to its popularity as a shipping route. Sample (a) in Figure S1 shows an example of such an error: from a data product standpoint, watercraft are not permanent fixtures and such artifacts should be excluded from the model; however, without manually editing the dataset or masking large water bodies according to an external dataset, it is impossible for the model to handle watercraft appropriately. This behavior is a byproduct of the mono-temporal nature of the input imagery: integrating multiple observations within a single year into the preprocessing pipeline or model inputs might address this problem, but doing so is impossible given the acquisition schedule of NAIP imagery.
Another systematic error is the appearance of sun glint in the imagery, leading to spurious misclassifications in the Gulf Coast region: sample (f) in Figure S1 gives an example of such misclassifications in the Mississippi Sound region. While quality control restricts the amount of allowable sun glint in NAIP imagery, it is impossible to account for and correct all instances of sun glint, leading to critical errors of omission when classifying water bodies. Given that NAIP is rarely used for analysis of water and there is no widely used protocol for sun glint correction in NAIP imagery, these artifacts cannot be removed. Including more samples of water bodies with high levels of sun glint in the training set is necessary to address this issue.
While the unclassified class provides little information from a data product perspective, its poor producer’s accuracy (32.29%) is potentially concerning. The most likely cause for this is the fact that most unclassified pixels lie directly adjacent to or in the middle of forested regions due to the presence of shadows. While an annotator may strictly classify small groups of dark pixels as unclassified, the model tended to classify these pixels based on the adjacent land cover. An example of this phenomenon can be observed in sample (b) in Figure S1, where small shadows in the forest are simply classified as belonging to the surrounding forest instead of the unclassified class, which is likely a correct classification given the context. Table 3 supports this conclusion, where it is evident that a large number of unclassified pixels were classified as tree canopy/woody vegetation. While this is sub-optimal from a data quality standpoint, as the unclassified class has a moderately high user’s accuracy (86.04%) and the class provides no useful information on surface phenomena, it is unlikely that this has a large impact on the overall suitability of this product for a given task.
S3 2016 data product assessment
| Predicted Class | |||||||||
| True Class | Open Water | Impervious Structures | Impervious Surfaces | Barren Land | Tree Canopy/Woody Vegetation | Low Vegetation/ Herbaceous | Cultivated Crops | Unclassified | Total |
| Open Water | 565 | 1 | 2 | 10 | 30 | 35 | 5 | 24 | 672 |
| Impervious Structures | 0 | 75 | 11 | 2 | 3 | 11 | 0 | 0 | 102 |
| Impervious Surfaces | 0 | 7 | 209 | 18 | 1 | 23 | 0 | 4 | 262 |
| Barren Land | 6 | 3 | 67 | 1,464 | 24 | 495 | 246 | 2 | 2,307 |
| Tree Canopy | 0 | 0 | 1 | 7 | 14,278 | 485 | 10 | 14 | 14,795 |
| Low Vegetation | 3 | 1 | 18 | 127 | 249 | 3,372 | 308 | 8 | 4,086 |
| Cultivated Crops | 0 | 0 | 1 | 43 | 0 | 101 | 793 | 0 | 938 |
| Unclassified | 6 | 7 | 11 | 6 | 1,239 | 130 | 0 | 439 | 1,838 |
| Total | 580 | 94 | 320 | 1,677 | 15,824 | 4,652 | 1,362 | 491 | 25,000 |
| Class Name | User’s Accuracy (%) | Producer’s Accuracy (%) | F1 Score (%) | IoU (%) | Cohen’s |
| Open Water | 97.41 | 84.08 | 90.26 | 82.24 | 0.9001 |
| Impervious Structures | 79.79 | 73.53 | 76.53 | 61.98 | 0.7644 |
| Impervious Surfaces | 65.31 | 79.77 | 71.82 | 56.03 | 0.7149 |
| Barren Land | 87.30 | 63.46 | 73.49 | 58.10 | 0.7126 |
| Tree Canopy/Woody Vegetation | 90.23 | 96.51 | 93.26 | 87.38 | 0.8265 |
| Low Vegetation/Herbaceous | 72.48 | 82.53 | 77.18 | 62.84 | 0.7237 |
| Cultivated Crops | 58.22 | 84.54 | 68.96 | 52.62 | 0.6751 |
| Unclassified | 89.41 | 23.88 | 37.70 | 23.23 | 0.3571 |
| Macro Average | 80.02 | 73.54 | 73.65 | 60.55 | 0.7093 |
| Overall | Accuracy: 84.78 | 73.96 | 0.7751 |
S4 Pre-training dataset visualization